OpenStudy (anonymous):

How do I solve this equation: 4500 = 1500(e)^(.04*x)

Nnesha (nnesha):

same like last one :-) hint: \[\huge\rm ln e^x = x \] natural log can cancel out e bec ln is inverse of e

OpenStudy (anonymous):

@Nnesha I divided by 1500 on both sides and got: 3 = e^(.04*x) So now I do: lne^(.04*x)?

Nnesha (nnesha):

yep that's right :-)

OpenStudy (anonymous):

@Nnesha I'm still confused on how I get the x out of the equation

Nnesha (nnesha):

so ln can cancel e \[\huge\rm \cancel{ln e}^x = x ~~~ \cancel {ln e}^p = p\]

OpenStudy (owlcoffee):

\[4500=1500e ^{0.4x}\] This is a typical equation that combines logarithms and exponents, it is indeed a good way to put in practice those little algebraic properties. Now, since the variable is in the exponent, we want to get rid of it, but first let's isolate the e^0.4x : \[e ^{0.4x}=\frac{ 4500 }{ 1500 }\] Now, in order to get that 0.4x to the numerator, we will take a natural logarithm on both sides: \[\ln(e ^{0.4x})=\ln (\frac{ 4500 }{ 1500 })\] and by logarithmic properties: \[0.4x(lne)=\ln (\frac{ 4500 }{ 1500 })\]

Nnesha (nnesha):

to get rid of e you should always apply "ln" log to get ride of "ln" you should always apply e

OpenStudy (anonymous):

@Nnesha Sorry I'm still very confused. If I now have: ln(3) = ln(e)^(.04x) then do I cross out the ln on both sides and have 3 = e^(.04x) ? I feel like I'm going in circles

Nnesha (nnesha):

okay "cross out the ln both sides" nope you don't have to cross ln then question would be same l 3 = e^(.04x) our purpose to take ln is to cancel out e and to bring the exponent down